At December 31, 2017, the balance sheet of Meca International included the following shareholders' equity accounts:

Shareholders’ Equity ($ in millions)
Common stock, 65 million shares at $1 par $ 65
Paid-in capital—excess of par 490
Retained earnings 600


Required:
Assuming that Meca International views its share buybacks as treasury stock, record the appropriate journal entry for each of the following transactions: (If no entry is required for a transaction/event, select "No journal entry required" in the first account field. Enter your answers in millions (i.e., 10,000,000 should be entered as 10).)

  1. On February 12, 2018, Meca reacquired 2 million common shares at $19 per share.
  2. On June 9, 2019, Meca reacquired 3 million common shares at $14 per share.
  3. On May 25, 2020, Meca sold 3 million treasury shares at $22 per share—determine cost as the weighted-average cost of treasury shares.
  4. For the previous transaction, assume Meca determines the cost of treasury shares by the FIFO method.
Answer #1
Meca International                              -  
Date Account Debit Credit Calculation
12-Feb Treasury Stock               38 =2*19
Cash               38
(To record 2 million common shares reacquired at $19 per share)
9-Jun Treasury Stock               42 =3*14
Cash               42
(To record 3 million common shares reacquired at $14 per share)
Weighted-average cost
25-May Cash               66 =3*22
Treasury Stock               48 =3*16
Paid in Capital from Treasury Stock               18
(To record sale of 3 million treasury shares at $22 per share)
FIFO cost
25-May Cash               66 =3*22
Treasury Stock               52 =(2*19)+(1*14)
Paid in Capital from Treasury Stock               14
(To record sale of 3 million treasury shares at $22 per share)
Workings
Date Weighted-average cost Shares Cost Total Cost
12-Feb Common shares reacquired                 2               19                             38
9-Jun Common shares reacquired                 3               14                             42
Totals                 5                             80
Weighted-average cost               16 =80 / 5